OpenStudy (anonymous):

Express cos^2 x in terms of cos2x

OpenStudy (usarmy3947):

cos (2x) = 1 - 2sin^2( x)

OpenStudy (anonymous):

are you getting it

OpenStudy (usarmy3947):

ah you there @jonathan283

zepdrix (zepdrix):

The Cosine Double Angle Formula comes in several forms, I think we want to use this one: \[\Large \color{royalblue}{\cos(2x)=2\cos^2x-1}\] Solve for cos^2x.